Alamo's customer service department follows up on customer complaints by telephone inquiry. During a recent period, the department initiated 15,700 calls and incurred costs of $329,700. If 3,150 of these calls were for the company's wholesale operation (the remainder were for the retail division), costs allocated to the retail division under activity-based costing system should amount to:
Answer #1

calculation of costs allocated to the retail division under activity based costing system:

(a) Total cost incurred =$ 329700

(b) Total number of calls initiated = 15700

(c) Number of calls related to wholesale operations = 3150

(d) Number of calls related to retail division = 12550 (15700-3150)

(e) cost allocated to the retail division = 329700*12550/15700= $ 263550.